# Learning the Geometric Mechanics of Robot Motion Using Gaussian Mixtures

### v1

@@ II-C Pre-processing to Produce a Ruled Surface
To enhance our algorithm’s ability to produce the desired linear relationship, we first construct a GMM with the sample data, and then reprocess the data to create extrapolations of the linear relationship, as follows.
@@
    Produce a GMM model $\{(w_{k},\mu_{k},S_{k})\}_{k=1}^{N}$ consisting of $N$ Gaussians, where the $\mu_{k}$ is the mean of the $k$-th constituent Gaussian, whose variance is $S_{k}$ and whose relative weight is $w_{k}$. For each of these Gaussians, compute $A_{k}$ such that $v_{b}={\mu_{v_{b}}}+A_{k}(\dot{r}-\mu_{\dot{r}})$ the TLS estimate of the motility map $A(\mu_{r})$.
@@
    Add the augmenting dataset to the original dataset and recompute the GMM to produce an “Augmented GMM”
@@
Assuming the Gaussians each have $c$ data points, this process creates an additional factor of $\alpha$ data points, each of which is the linear combination of $c$ points extrapolated by a factor of $\beta$. If the original data was affine in its relationship between $\dot{r}$ and $v_{b}$, the augmentation with a $\beta>1$ extrapolates this relationship to a larger distance along the $\dot{r}$ directions, which would result in an Augmented GMM that respects this affine relationship over a larger range of $\dot{r}$.
Sensible choices for the meta-parameters require choosing $c$ large enough to sufficiently reduce the noise of the affine model by averaging nearby points so that an extrapolation by factor $\beta$ will still be accurate enough.
For our example here we chose the number of Gaussians as $60$, $\alpha=1,\beta=1.2,c=5$.
II-D Model types
Since our goal is to improve the ability to construct data-driven motility map models, we chose to compare robot motion tracking data (“ground truth”) to our new GMM based approach, and to two additional model types derived from the approach of [[bittner2018geometrically]] and [[zhao2022walking]]. We challenged our models with extrapolation, similar to [[zhao2022walking]], by building a model using training data from an intermediate turning gait, and testing against both sharper and more moderate turns. This is because only models that can correctly extrapolate motions outside of training data are usable for the optimization framework shown in [[bittner2018geometrically]] and [[bittner2021optimizing]].
For each training dataset, we computed a motility map model following the construction in [[bittner2018geometrically]], namely first separating the data into bins by “phase”, i.e. where the robot was within its cycle of motion. For each bin we computed the average body velocity, and built a first order Taylor series expansion of the motility map around this phase-averaged behavior. Finally, we interpolated the bin averages and the tensors representing the Taylor series using a Fourier series to produce a body velocity estimation function of the form $v_{b}(\phi,r,\dot{r}):=v_{0}(\phi)+\Delta(\phi,r)\dot{r}$ giving body velocity as a function of phase, shape, and shape velocity. Here the average body velocity as a function of phase was $v_{0}(\phi)$, and we added a first order correction $\Delta(\phi,\cdot)\cdot$.
The two additional model types we compared with are the “Geometric” model given by $v_{b}(\phi,r,\dot{r})$, and the “Phase” model given by $v_{0}(\phi)$ which does not utilize any information about how the shape changes away from the average training shape at that phase.
II-E Loss Function
To select a good set of meta-parameters, we chose a loss function representing a non-dimensionalized modeling error. For each stride, i.e. full gait cycle, we consider the forward (initial body frame $x$), sideways (initial body frame $y$), and heading (initial body frame $\theta$) motion produced. Ignoring any possible correlation of these variables, we took the loss to be the total $z$-score of the $x$, $y$, and $\theta$ predictions with respect to the corresponding variable in the ground truth data. This means that with our choice of loss function, the loss of a random permutation of the ground truth data against itself was set to approach $3$, and the loss of a prefect prediction is zero.
III Results
Refer to captionRefer to captionRefer to caption
@@ III Results
Figure 2: Comparison of model extrapolation to less and more turning. We compared the predictions of the models for a recording of $6$ strides (cycles) of motion, after training on data from an intermediate turning rate (steering parameter $s=0.50$ of [[zhao2020multi]]). We attempted to predict both noticeably lesser ($s=0.25$; left) and noticeably greater ($s=0.75$; middle) turning rates; the different turning radii are demonstrated by a circle fitted to the motion tracking data (dashed green line and circular markers for centers; right plot; reproduced from Fig. 3.9 of [[Zhao-2021-PhD]]). In each case we plotted the ground truth motion capture body centroid (thick red lines, triangle markers), the Geometric model (blue lines, circle markers), the Phase model (green lines, circle markers), and the new GMM model (dark gold lines, diamond markers) We plotted the markers at the same phase in each cycle of motion, to make it easier to understand how much the robot moved in a cycle. Results demonstrate that while the Geometric model is noticeably better than the Phase model, the GMM model is noticeably superior to both.
@@
To get a general sense of the difference between the three model types, we provided Fig. 2 which shows two examples of observed body centroid trajectories and the predicted motion based on the various models.
@@
We obtained better insight into the predictive ability of these models with Fig. 3, which shows the motion after a full cycle – either observed or predicted – and also the prediction errors in comparison with the variability in the ground truth data. We plotted this as a kernel smoothed density plot to give a (non-parametric) sense of how well these predicted distributions overlap.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Figure 3: Distribution of final positions after a cycle of motion. We plot a kernel smoothed density estimate (KSDE) of the forward ($x$), sideways ($y$), and turn ($\theta$) motion produced by each stride (i.e. full gait cycle), and their error distributions. To provide a reference variance allowing the relative size of the modeling error to be understood by the reader, we plotted the ground truth data minus its mean in the error distributions. We plotted ground truth (green), the [[bittner2018geometrically]] gait-based geometric model prediction (orange), and the GMM model prediction (purple). We trained the models on intermediate turning data (“turn parameter” $s=0.50$) and plotted the extrapolated results for a slower turn ($s=0.25$; $x$, $y$ top row, $\theta$ left) and a faster turn ($s=0.75$; $x$, $y$ middle row, $\theta$ right). For the $\theta$ distribution we plotted the observed values (vertical tics at the bottom of the plots), the distribution medians (thick vertical lines), and a scale bar for the width of the smoothing kernel (thick black line).
The results show that the GMM error distribution is much closer to zero in every coordinate, and contains zero error close to its mode. The variability in GMM model prediction is of a similar order to the variability in the data itself. Because the BigANT robot was commanded to produce identical motions each cycle, this suggests the possibility that the model variability may be due to system noise that cannot be captured by the variables we measure, rather than modeling error.
IV Discussion
@@
Our results show that using a GMM to model the motility map produced predictions that were substantially more accurate than the previous data-driven geometric mechanics modeling approaches cited. We have shown this for an extrapolation problem — for motions substantially different from the training data (see Fig. 2 right), suggesting that these models can be useful for planning and for iterative gait optimization similar to that shown in [[bittner2018geometrically, bittner2021optimizing]].
@@
Obviously, the results here are preliminary and much work remains. Our modeling approach needs to be tested with a variety of datasets, including testing with data that contain some hysteresis effects, and data that are less linear in the $\dot{r}$ dependence of the motility map. It would be natural to extend the models to the Shape Underactuated Dissipative Systems of [[bittner2022data]], and use methods for online update of GMM-s to produce an online adaptive version of our method. Another natural direction is to automate and improve the selection of meta-parameters, and to try other established manifold learning approaches for constructing models.
Acknowledgments: funding for this work was provided by NSF CPS 2038432; dataset was collected with funding from ARO MURI W911NF-17-1-0306 using equipment funded by ARO DURIP W911NF-17-1-0243.
[…]

### v2

@@ 2.3 Pre-processing to Produce a Ruled Surface
To enhance our algorithm’s ability to produce the desired linear relationship from small samples, we first construct a GBR with the sample data, and then reprocess the data to create extrapolations of the linear relationship, as follows.
@@
    Produce a GBR model $\{(w_{k},\mu_{k},S_{k})\}_{k=1}^{N}$ consisting of $N$ Gaussians, where the $\mu_{k}$ is the mean of the $k$-th constituent Gaussian, whose variance is $S_{k}$ and whose relative weight is $w_{k}$. For each of these Gaussians, compute $A_{k}$ such that $v_{b}={\mu_{v_{b}}}+A_{k}(\dot{r}-\mu_{\dot{r}})$ the TLS estimate of the motility map $A(\mu_{r})$.
@@
    Add the augmenting dataset to the original dataset and recompute the GBR to produce an “Augmented GBR (A-GBR)”
@@
Assuming the Gaussians each have $c$ data points, by using this process we created an additional factor of $\alpha$ data points, each of which is the linear combination of $c$ points extrapolated by a factor of $\beta$. If the original data was affine in its relationship between $\dot{r}$ and $v_{b}$, by augmenting it with a $\beta>1$ we extrapolated this relationship to a larger distance along the $\dot{r}$ directions. By this we produced an A-GBR that respects this affine relationship over a larger range of $\dot{r}$ than a naive GBR model.
We needed to choose $\alpha$, $\beta$, and $c$. We chose $c$ large enough to sufficiently reduce the noise of the affine model by averaging nearby points so that our extrapolation by factor $\beta$ seemed to provide a benefit. We chose $\alpha$ arbitrarily, but small enough not to noticeably slow the computation time. The parameters we used are: number of Gaussians $60$, $\alpha=1$, $\beta=1.2$, $c=5$.
Figure 2: BigAnt Robot.The BigAnt robot has six 1-DoF legs each using a 4-bar linkage which was carefully designed balancing leg clearance and the changes in gearing ratio (Zhao, 2021, ch. 2). It was designed as a demonstration of the benefits of the “Plates and Reinforced Flexures” approach to robot design and fabrication (Fitzner et al., 2017): its chassis uses only two materials (foam core board and fiber reinforced tape), costs about \$US 20, and takes a competent student about one day to make. We have demonstrated the ability of this robot to go over rough terrain [Wave Field Video] and through snow [BigAnt in Snow Video].
Figure 3: Multipod Robot. The Multipod robots are a collection of segmented modular robots with spring steel legs, going from 4 legs (2 pairs) through to 12 legs (6 pairs). These robots were used to study how locmotion dynamics changes when the number of legs increase, leading in part to the results of Zhao et al. (2022).
@@
Figure 4: Comparison of model extrapolation to less and more turning. We compared the predictions of the models for a recording of $6$ strides (cycles) of motion, after training on data from an intermediate turning rate (steering parameter $s=0.50$ of Zhao and Revzen (2020)). We attempted to predict both noticeably lesser ($s=0.25$; top) and noticeably greater ($s=0.75$; bottom) turning rates;. In each case we plotted the ground truth motion capture body centroid (thick green lines, triangle markers), the Geometric model (orange lines, circle markers), the Phase model (pink lines, circle markers), and the A-GBR model (purple lines, diamond markers) We plotted the markers at the same phase in each cycle of motion, to make it easier to understand how much the robot moved in a cycle. Results demonstrate that while the Geometric model is noticeably better than the Phase model, A-GBR is noticeably superior to both.
Figure 5: A-GBR performance on the Multipod. We trained A-GBR with data from intermediate phase offset gaits ($\Delta\phi=[1.05\pi,1.1\pi,1.15\pi]$) and tested its performance on gaits with both lower and higher $\Delta\phi$. We provide similar trajectory plot as in Fig. 4 (top plots $\Delta\phi=[0.85\pi,0.9\pi,0.95\pi,0.97\pi,\pi]$; bottom plots $\Delta\phi=[1.2\pi,1.25\pi,1.3\pi,1.35\pi,1.4\pi]$). While the Geometric model is noticeably better than the Phase model, A-GBR is noticeably superior to both.
2.4 Model types
Since our goal is to improve the ability to construct data-driven motility map models, we chose to compare robot motion tracking data (“ground truth”) to our A-GBR estimates, and compared these estimates to estimates obtained from the approach of Bittner et al. (2018) and Zhao et al. (2022). We challenged our models with extrapolation, similar to Zhao et al. (2022), by building a model using training data from an intermediate turning gait, and testing against both sharper and more moderate turns. The ability to extrapolate well is critical for making a model usable for the optimization framework shown in Bittner et al. (2018) and Bittner and Revzen (2021).
2.4.1 Geometric Model
For each training dataset, we computed a motility map model following the construction in Bittner et al. (2018), namely first separating the data into bins by “phase”, i.e. where the robot was within its cycle of motion. For each bin we computed the average body velocity, and built a first order Taylor series expansion of the motility map around this phase-averaged behavior. Finally, we interpolated the bin averages and the tensors representing the Taylor series using a Fourier series to produce a body velocity estimation function of the form $v_{b}(\phi,r,\dot{r}):=v_{0}(\phi)+\Delta(\phi,r)\dot{r}$ giving body velocity as a function of phase, shape, and shape velocity. Here the average body velocity as a function of phase was $v_{0}(\phi)$, and we added a first order correction $\Delta(\phi,\cdot)\cdot$.
2.4.2 Phase Model
The “Phase” model uses only the the phase-averaged $v_{b}$, i.e. is equivalent to the “Geometric” model with $\Delta:=0$.
2.5 Loss Function
To select a good set of meta-parameters, we chose a loss function representing a non-dimensionalized modeling error. For each stride, i.e. full gait cycle, we consider the forward (initial body frame $x$), sideways (initial body frame $y$), and heading (initial body frame $\theta$) motion produced. Ignoring any possible correlation of these variables, we took the loss to be the total $z$-score of the $x$, $y$, and $\theta$ predictions with respect to the corresponding variable in the ground truth data. This means that with our choice of loss function, the loss of a random permutation of the ground truth data against itself was set to approach $3$, and the loss of a perfect prediction is zero.
3 Results
To evaluate the performance of our model, we present results using motion capture data from the BigANT (Fig. 2) BIRDS-Lab (2021a) and Multipod (Fig. 3) BIRDS-Lab (2021b) robots.
Data for both BigANT and Multipod were collected using a Qualisys motion capture system consisting of 10 Oqus-310+ cameras operating at 120 fps as described in BIRDS-Lab (2021a, b). We trained the Phase, Geometric, and A-GBR models using shape space data represented by motion tracking marker locations relative to a body frame. The body frame we used in each case was the body frame available in the publicly available dataset; optimal selection of the body frame, as in Hatton et al. (2011), might improve on our results.
3.1 BigANT Results
BigANT is a hexapod robot with one degree of freedom per leg. The analyzed gaits induce BigANT to steer in varying rates, as defined by Zhao and Revzen (2020). In this study, we assess the extrapolation capability of A-GBR by training it on data with an intermediate steering rate and testing it on both lower and higher steering rate recordings.
@@ 3.1 BigANT Results
To get a general sense of the difference between the three model types when applied to BigANT, we provided Fig. 6 and Fig. 4 which shows examples of observed body centroid trajectories and the predicted motion based on the various models.
Figure 6: Turning rate demonstration. The different turning radii are demonstrated by a circle fitted to the motion tracking data (dashed green line and circular markers for centers; right plot; reproduced from Fig. 3.9 of Zhao (2021)).
3.2 Multipod Results
The Multipod robots are a family of serpentine multi-legged robots varying from 4 to 12 legs. Each leg is (functionally) movable in 2 DoF, and the robots move by commanding the legs in a traveling wave defined by the phase offset between adjacent segments. In this study, we focus on the YRY (yaw-roll-yaw) configuration 6-legged Multipod, as defined by Zhao and Revzen (2020); Zhao (2021). We trained A-GBR on gaits intermediate phase offsets ($\Delta\phi=[1.05\pi,1.1\pi,1.15\pi]$) and tested on both lower ($\Delta\phi=[0.85\pi,0.9\pi,0.95\pi,0.97\pi,\pi]$) and higher ($\Delta\phi=[1.2\pi,1.25\pi,1.3\pi,1.35\pi,1.4\pi]$) phase offsets, as defined by Zhao (2021). Figure 5 illustrates the robot body frame trajectories corresponding to these different gaits.
3.3 Prediction and Residual Error Distributions
@@
We obtained better insight into the predictive ability of these models (BigANT Fig. 7 and Fig. 8; Multipod Fig. 9 and Fig. 10) by considering the distributions of the predictions of each model and the distributions of the residual errors of the models plotted on the backdrop of the distribution of the ground truth against its mean for scale. We plotted these as kernel-smoothed density plots to give a (non-parametric) sense of how well these predicted distributions overlap.
@@
Figure 7: Distribution of final positions after a cycle of BigANT motion. We ploted a kernel-smoothed density estimate (KSDE) of the forward ($x$), sideways ($y$), and turn ($\theta$) motion produced by each stride (i.e. full gait cycle). We plotted ground truth (green), the Bittner et al. (2018) gait-based geometric model prediction (orange), and the GMM model prediction (purple). We trained the models on intermediate turning data (“turn parameter” $s=0.50$) and plotted the extrapolated results for a slower turn ($s=0.25$; $x$, $y$, $\theta$ left) and a faster turn ($s=0.75$; $x$, $y$, $\theta$ right). For the $\theta$ distribution we plotted the observed values (vertical tics at the bottom of the plots), the distribution medians (thick vertical lines), and a scale bar for the width of the smoothing kernel (thick black line).
Figure 8: Distribution of the prediction residual of final positions after a cycle of BigANT motion. We ploted a KSDE of residual error distributions of the forward ($x$), sideways ($y$), and turn ($\theta$) motion produced by each stride (i.e. full gait cycle). To provide a reference variance allowing the relative size of the modeling error to be understood by the reader, we plotted the ground truth data minus its mean in the error distributions. The model and training setting are the same as in Fig.7. For the $\theta$ distribution we also plotted the observed values, the distribution medians, and a scale bar for the width of the smoothing kernel as in Fig.7.
Figure 9: Distribution of final positions after a cycle of Multipod motions. We plot a KSDE of the forward ($x$), sideways ($y$), and turn ($\theta$) motion produced by each stride (i.e. full gait cycle). The model setting is the same as in Fig. 7. We trained the models on intermediate phase offset ($\Delta\phi=[1.05\pi,1.1\pi,1.15\pi]$) and plotted the extrapolated results for lower phase offsets ($\Delta\phi=[0.85\pi,0.9\pi,0.95\pi,0.97\pi,\pi]$, left) and higher phase offsets ($\Delta\phi=[1.2\pi,1.25\pi,1.3\pi,1.35\pi,1.4\pi]$, right). For the $\theta$ distribution, we present both the predicted values and the corresponding residuals.
Figure 10: Distribution of the prediction residual of final positions after a cycle of Multipod motions. We plot a KSDE of of residual error distributions of the forward ($x$), sideways ($y$), and turn ($\theta$) motion produced by each stride (i.e. full gait cycle) The model and training setting are the same as in Fig. 9. For the $\theta$ distribution, we present 1-D KSDE of the residuals as described in Fig. 8.
4 Discussion
@@
Our results show that using A-GBR to model the motility map produced predictions that were substantially more accurate than the previous data-driven geometric mechanics modeling approaches cited. We have shown this for an extrapolation problem — for motions substantially different from the training data (see Fig. 4, Fig. 5 and Fig. 6), suggesting that these models can be useful for planning and for iterative gait optimization similar to that shown in Bittner et al. (2018); Bittner and Revzen (2021).
@@
Obviously, the results here are preliminary and much work remains. Our modeling approach needs to be tested with a variety of datasets, including testing with data that contain notable hysteresis effects, and data that are less linear in the $\dot{r}$ dependence of the motility map. It would be natural to extend the models to the Shape Underactuated Dissipative Systems of Bittner et al. (2022), and use methods for online update of GMM-s to produce an online adaptive version of A-GBR. Another natural direction is to automate and improve the selection of meta-parameters, and to try other established manifold learning approaches for constructing models.
Funding
Funding for this work was provided by NSF CPS 2038432; dataset was collected with funding from ARO MURI W911NF-17-1-0306 using equipment funded by ARO DURIP W911NF-17-1-0243.